# RECAST: Strengthening LLMs’ Complex Instruction Following with Constraint-Verifiable Data

### v2

@@ Abstract
Large language models (LLMs) are increasingly expected to tackle complex tasks, driven by their expanding applications and users’ growing proficiency in crafting sophisticated prompts. However, as the number of explicitly stated requirements increases (particularly more than $10$ constraints), LLMs often struggle to accurately follow such complex instructions. To address this challenge, we propose RECAST, a novel framework for synthesizing datasets where each example incorporates far more constraints than those in existing benchmarks. These constraints are extracted from real-world prompt-response pairs to ensure practical relevance. RECAST enables automatic verification of constraint satisfaction via rule-based validators for quantitative constraints and LLM-based validators for qualitative ones. Using this framework, we construct RECAST-$30$K, a large-scale, high-quality dataset comprising $30$k instances spanning $15$ constraint types. Experimental results demonstrate that models fine-tuned on RECAST-$30$K show substantial improvements in following complex instructions. Moreover, the verifiability provided by RECAST enables the design of reward functions for reinforcement learning, which further boosts model performance on complex and challenging tasks.
Figure 1: Challenges in complex instruction following and improved performance via RESAST-30k. Left: Real-world examples illustrating how LLMs fail to follow complex instructions. Top-right: Performance degradation of LLMs as the number of constraints increases. Bottom-right: Comparison of instruction-following performance between models fine-tuned on RECAST and their corresponding Instruct variants (e.g., Qwen2.5-7B-Instruct).
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in solving various NLP tasks [1, 2, 3, 4, 5, 6, 7] and are widely used in practical scenarios [8, 9, 10, 11]. However, they still struggle with complex multi-constraint scenarios that are prevalent in real-world applications, such as constitutional AI systems requiring adherence to multiple principles simultaneously [12, 13, 14, 15], and enterprise assistants managing detailed business rules [16, 17, 18]. As shown in Figure 1, when instruction complexity increases, with multiple explicit constraints in a single prompt, even advanced models like GPT-4o [19] show marked performance degradation, which has been emphasized in multiple benchmarks [20, 21, 22, 23].
While recent research efforts have focused on improving instruction-following capabilities through automated data generation [24, 25, 26, 27, 28, 29, 30, 31, 32] and improved training methodologies[33, 34, 35, 27, 29, 36, 37], several challenges remain underexplored. First, existing instruction datasets primarily consist of examples with simple constraints [27, 30, 31] that inadequately reflect the richness and complexity of real-world queries. Second, conventional training methodologies prioritize overall response quality [35, 27] over fine-grained constraint adherence, leaving models prone to violating specific requirements when confronted with intricate, multi-constraint instructions.
@@
Based on RECAST, we constructed a high-quality, diverse, and constraint-verifiable dataset, RECAST-30K, which encompasses 15 common types of constraints. Using only 30K data, the base model, such as Qwen2.5-7B-Base [38] trained on RECAST-30K outperforms the corresponding instruct model, which was trained on a larger-scale dataset. The RLVC approach further enhances constraint compliance, yielding substantial improvements in multi-constraint satisfaction. Through detailed analysis, we identify key factors affecting instruction-following reliability, finding that both constraint types and quantity impact model performance in complex scenarios. To summarize the highlights of our study:
@@ 2 Related Work
Recent research has explored various approaches to automatically construct complex instruction datasets. WizardLM [31] introduced Evol-Instruct, which iteratively increases instruction difficulty through rewriting. Conifer [29] employs a top-down approach using GPT-4 to generate multi-level constraint instructions. ULTRAIF [39] decomposes instructions into basic queries and constraints before recombination, while AIR [25] applies an iterative refinement framework that progressively adds constraints through a model-judge cycle. Other approaches include MUFFIN [40], which builds multi-faceted tasks by expanding input features, and CRaB [28], which infers new constraints from existing instruction-response pairs through back-translation. Unlike these methods, our RECAST approach uniquely focuses on collecting diverse constraint types from high-quality responses and incorporates explicit verification mechanisms for both rule-based and model-based constraints, addressing limitations in existing methods regarding constraint complexity, diversity, and quality assurance.
@@
Recent algorithmic innovations have significantly advanced instruction following capabilities. AutoIF [34] implements a self-dialogue generation process with execution-based verification, where models produce instructions, verification code, and test cases, filtering training data through executable feedback. RNR [37] extracts roles and rules from existing instructions to generate rule-compliant responses, while SPaR [33] employs self-play tree search refinement through multiple rounds of model self-adversarial interaction to optimize response quality. The approach in [35] utilizes discriminative generation, applying a discriminator model to filter generated samples for higher-quality supervision. Unlike these methods, our RLVC framework specifically leverages constraint verifiability, designing multi-channel reward signals that treat each constraint as an independent optimization objective. This fine-grained reinforcement learning approach enables more precise feedback on constraint satisfaction, substantially improving both the accuracy and consistency of complex instruction following.
@@ 4.3 Policy Optimization
We employ Group Relative Policy Optimization (GRPO) [41] as our policy learning algorithm. GRPO generates multiple candidate responses for the same instruction and computes advantage estimates through intra-group comparisons, allowing the model to learn the relative quality differences among various responses to the same instruction.
@@ 5.1 Setups
$\text{HSR}=\frac{1}{|D|}\sum_{(x,y)\in D}\frac{1}{|C(x)|}\prod_{c_{i}\in C(x)}f(x,y,c_{i})$ (5)
@@
For the FollowBench evaluation, we similarly adopt the Hard Constraint Satisfaction Rate defined in the original benchmark. For all LLM-based evaluations, unless otherwise specified, we default to using GPT-4o [19] for assessment.
@@
For all experiments, we choose Qwen2.5-7B [38] and Llama-3.1-8B [42] as the base models. The same experimental setup is used for SFT across all datasets. We use the model after SFT as the base model of RLVC for training. Specific experimental details can be found in Appendix C.
@@ 5.2.1 Evaluation on RECAST-Test
A consistent performance degradation is observed across all models as constraint complexity increases from Level 1 to Level 4. This systematic decline highlights the fundamental challenge LLMs face when processing and satisfying multiple constraints simultaneously. Notably, rule-based constraints present greater difficulties than model-based constraints across all model classes. Among tested models, Gemini-2.5-Pro demonstrates superior constraint handling capabilities with a $39.75$% average satisfaction rate.
@@
Models fine-tuned on RECAST-30K demonstrate substantial improvements in constraint satisfaction capabilities. The RECAST-30K-SFT variant achieves a $31.25$% average satisfaction rate on Qwen-2.5-7B, significantly outperforming both instruction-tuned models and alternative fine-tuning approaches. This enhancement underscores the effectiveness of our constraint-focused dataset in developing robust instruction-following capabilities, particularly for navigating complex, multi-constraint scenarios.
@@
The application of RLVC further enhances model performance across all difficulty levels. RECAST-30K-RLVC variants consistently outperform their SFT counterparts, achieving a $32.33$% average satisfaction rate on Qwen-2.5-7B. This improvement is particularly pronounced in higher difficulty levels (Levels 3-4), where constraint interactions become more complex. These gains confirm that our constraint-specific reinforcement learning approach effectively optimizes for simultaneous satisfaction of multiple requirements, addressing a key challenge in practical instruction following.
@@ 5.3 Ablations and Analysis
To isolate component contributions, we evaluated RECAST variants with specific elements removed. Complete RECAST-30k consistently outperforms all ablated configurations. Instruction-Only Enhancement exhibits the most significant performance degradation ($12.21$% and $17.37$% decreases for Llama and Qwen), revealing that constraint-augmented responses without corresponding response modifications create instruction-response inconsistencies that confuse the model, substantially hampering performance. Response-Only Enhancement performs comparably to Without RECAST, suggesting that constraint specification in instructions drives most performance gains. These results confirm that RECAST’s complete instruction-response synthesis pipeline is essential for optimal constraint satisfaction capabilities.
@@
To analyze the effectiveness of our reinforcement learning approach, we tracked both reward metrics and constraint satisfaction performance throughout the RLVC training process. Figure 5 illustrates the reward progression. This reward optimization corresponds directly to improved constraint satisfaction capabilities. As shown in the evaluation data, the Rule-based Constraint Satisfaction Rate steadily improves from an initial $5$% at step 00 to $11$% at step $200$. The reward growth (approximately $5$% increase) reflects the inherent difficulty of the constraint satisfaction task. Unlike simpler RL objectives with clearer optimization landscapes, constraint satisfaction requires balancing multiple, potentially competing requirements simultaneously. This complexity creates a challenging reward space where even small improvements represent meaningful learning. These results confirm that our constraint-specific reward design effectively guides model optimization toward improved rule-based constraint handling capabilities, demonstrating the value of targeted reinforcement learning for complex instruction-following tasks.
@@ C.1 Baseline
  • Conifer[29]: Constructs a GPT-4-refined dataset with hierarchical constraint instructions and employs curriculum tuning to enhance complex instruction-following.
  • Crab[28]: Enhances LLMs by back-translating model responses into constraint-rich instructions using a stronger teacher model.
  • I-SHEEP[43]: Aligns LLMs from scratch via iterative self-generated instruction–response tuning without external supervision.
  • MUFFIN[40]: Diversifies instruction data by scaling input facets per task, yielding more robust multi-faceted alignment.
@@
  • Evol-Instruct[31]: Generates complex instructions through iterative rewriting, producing progressively harder instruction–response pairs.
  • Suri[27]: Back-translates long-form texts into multi-constraint prompts, improving structured generation over extended contexts.
  • Tülu 3 Persona IF[44]: Synthesizes constraint-following data from diverse persona prompts to enhance controllable instruction tuning.
@@ C.2 General Capability Evaluation Benchmark
  • FollowBench [21] comprehensively evaluates instruction-following capability through five types of fine-grained constraints (Content, Situation, Style, Format, and Example) using a multi-level mechanism that incrementally adds constraints to initial instructions. This benchmark systematically measures models’ ability to satisfy multiple constraints simultaneously, with difficulty levels ranging from simple single-constraint instructions to complex multi-constraint scenarios. The evaluation employs metrics including Hard Satisfaction Rate (HSR) for complete constraint satisfaction and Soft Satisfaction Rate (SSR) for partial compliance.
  • IFEval [45] assesses models’ ability to follow explicit formatting instructions through strict, verifiable metrics rather than subjective content evaluation. The benchmark focuses on precise adherence to specific requirements such as keyword inclusion, format specifications, and structural constraints. This allows for rigorous, code-based evaluation that eliminates ambiguity in measuring instruction compliance.
  • BBH [46] comprises 23 carefully selected challenging tasks from the BigBench dataset that test diverse cognitive capabilities including multistep arithmetic, algorithmic reasoning, language understanding, and world knowledge. These tasks are chosen for their high difficulty, objective metrics, and strong correlation with human preferences. The benchmark provides insights into models’ reasoning abilities across domains ranging from boolean expression evaluation to sarcasm detection.
  • GPQA [47] features expert-crafted questions from PhD-level domain specialists in biology, physics, and chemistry designed to be challenging for non-experts yet accessible to specialists. The dataset undergoes rigorous validation to ensure both factual accuracy and appropriate difficulty levels. Access restrictions and gating mechanisms protect against data contamination, maintaining the benchmark’s integrity for evaluating advanced scientific knowledge.
  • MuSR [48] tests models’ ability to integrate reasoning with long-context understanding through algorithmically generated complex problems averaging 1,000 words in length. The benchmark includes murder mysteries, object placement puzzles, and team allocation optimization tasks that require maintaining coherent reasoning across extended contexts. Few models achieve above-random performance, making it an effective discriminator of advanced reasoning capabilities.
  • MMLU-PRO [49] represents an enhanced version of the original MMLU benchmark, addressing previous limitations through increased difficulty and improved quality control. The benchmark presents 10 answer choices instead of 4, requires deeper reasoning, and underwent expert review to eliminate noisy or unanswerable questions. This refinement results in a more challenging and reliable assessment of multitask language understanding across diverse knowledge domains.

### v3

@@ Abstract
Large language models (LLMs) are increasingly expected to tackle complex tasks, driven by their expanding applications and users’ growing proficiency in crafting sophisticated prompts. However, as the number of explicitly stated requirements increases (particularly more than $10$ constraints), LLMs often struggle to accurately follow such complex instructions. To address this challenge, we propose RECAST, a novel framework for synthesizing datasets where each example incorporates far more constraints than those in existing benchmarks. These constraints are extracted from real-world prompt-response pairs to ensure practical relevance. RECAST enables automatic verification of constraint satisfaction via rule-based validators for quantitative constraints and LLM-based validators for qualitative ones. Using this framework, we construct RECAST-3030K, a large-scale, high-quality dataset comprising 3030k instances spanning $15$ constraint types. Experimental results demonstrate that models fine-tuned on RECAST-3030K show substantial improvements in following complex instructions. Moreover, the verifiability provided by RECAST enables the design of reward functions for reinforcement learning, which further boosts model performance on complex and challenging tasks.
Figure 1: Challenges in complex instruction following and improved performance via RECAST-30k. Left: Real-world examples illustrating how LLMs fail to follow complex instructions. Top-right: Performance degradation of LLMs as the number of constraints increases. Bottom-right: Comparison of instruction-following performance between models fine-tuned on RECAST and their corresponding Instruct variants (e.g., Qwen2.5-7B-Instruct).
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in solving various NLP tasks Ahn et al. (2024); Chakrabarty et al. (2024); Gómez-Rodríguez and Williams (2023); Huang and Chang (2023); Jiang et al. (2024a); Zhao et al. (2023); Zheng et al. (2023) and are widely used in practical scenarios Gu et al. (2024); Liu et al. (2024a); Sun (2023); Thirunavukarasu et al. (2023). However, they still struggle with complex multi-constraint scenarios that are prevalent in real-world applications, such as constitutional AI systems requiring adherence to multiple principles simultaneously Bai et al. (2022); Findeis et al. (2024); Kundu et al. (2023); Kyrychenko et al. (2025), and enterprise assistants managing detailed business rules De Michele et al. (2025); Grohs et al. (2023); Robino (2025). As shown in Figure 1, when instruction complexity increases, with multiple explicit constraints in a single prompt, even advanced models like GPT-4o Achiam et al. (2023) show marked performance degradation, which has been emphasized in multiple benchmarks Ferraz et al. (2024); Jiang et al. (2024b); Qin et al. (2024); Wen et al. (2024).
While recent research efforts have focused on improving instruction-following capabilities through automated data generation Liang et al. (2024a); Liu et al. (2025, 2024b); Pham et al. (2024); Qi et al. (2024); Sun et al. (2024); Wang et al. (2023); Xu et al. (2024); Zhou et al. (2024) and improved training methodologiesCheng et al. (2024); Dong et al. (2024); He et al. (2024); Pham et al. (2024); Sun et al. (2024); Wallace et al. (2024); Wang et al. (2024a), several challenges remain underexplored. First, existing instruction datasets primarily consist of examples with simple constraints Pham et al. (2024); Wang et al. (2023); Xu et al. (2024) that inadequately reflect the richness and complexity of real-world queries. Second, conventional training methodologies prioritize overall response quality He et al. (2024); Pham et al. (2024) over fine-grained constraint adherence, leaving models prone to violating specific requirements when confronted with intricate, multi-constraint instructions.
@@
Based on RECAST, we constructed a high-quality, diverse, and constraint-verifiable dataset, RECAST-30K, which encompasses 15 common types of constraints. Using only 30K data, the base model, such as Qwen2.5-7B-Base Yang et al. (2024) trained on RECAST-30K outperforms the corresponding instruct model, which was trained on a larger-scale dataset. The RLVC approach further enhances constraint compliance, yielding substantial improvements in multi-constraint satisfaction. Through detailed analysis, we identify key factors affecting instruction-following reliability, finding that both constraint types and quantity impact model performance in complex scenarios. To summarize the highlights of our study:
@@ 2 Related Work
Recent research has explored various approaches to automatically construct complex instruction datasets. WizardLM Xu et al. (2024) introduced Evol-Instruct, which iteratively increases instruction difficulty through rewriting. Conifer Sun et al. (2024) employs a top-down approach using GPT-4 to generate multi-level constraint instructions. ULTRAIF An et al. (2025) decomposes instructions into basic queries and constraints before recombination, while AIR Liu et al. (2025) applies an iterative refinement framework that progressively adds constraints through a model-judge cycle. Other approaches include MUFFIN Lou et al. (2023), which builds multi-faceted tasks by expanding input features, and CRaB Qi et al. (2024), which infers new constraints from existing instruction-response pairs through back-translation. Unlike these methods, our RECAST approach uniquely focuses on collecting diverse constraint types from high-quality responses and incorporates explicit verification mechanisms for both rule-based and model-based constraints, addressing limitations in existing methods regarding constraint complexity, diversity, and quality assurance.
@@
Recent algorithmic innovations have significantly advanced instruction following capabilities. AutoIF Dong et al. (2024) implements a self-dialogue generation process with execution-based verification, where models produce instructions, verification code, and test cases, filtering training data through executable feedback. RNR Wang et al. (2024a) extracts roles and rules from existing instructions to generate rule-compliant responses, while SPaR Cheng et al. (2024) employs self-play tree search refinement through multiple rounds of model self-adversarial interaction to optimize response quality. The approach in He et al. (2024) utilizes discriminative generation, applying a discriminator model to filter generated samples for higher-quality supervision. Unlike these methods, our RLVC framework specifically leverages constraint verifiability, designing multi-channel reward signals that treat each constraint as an independent optimization objective. This fine-grained reinforcement learning approach enables more precise feedback on constraint satisfaction, substantially improving both the accuracy and consistency of complex instruction following.
@@ 4.3 Policy Optimization
We employ Group Relative Policy Optimization (GRPO) Shao et al. (2024) as our policy learning algorithm. GRPO generates multiple candidate responses for the same instruction and computes advantage estimates through intra-group comparisons, allowing the model to learn the relative quality differences among various responses to the same instruction.
@@ 5.1 Setups
$\text{HSR}=\frac{1}{|D|}\sum_{(x,y)\in D}\prod_{c_{i}\in C(x)}f(x,y,c_{i})$ (5)
@@
For the FollowBench evaluation, we similarly adopt the Hard Constraint Satisfaction Rate defined in the original benchmark. For all LLM-based evaluations, unless otherwise specified, we default to using GPT-4o Achiam et al. (2023) for assessment.
@@
For all experiments, we choose Qwen2.5-7B Yang et al. (2024) and Llama-3.1-8B Grattafiori et al. (2024) as the base models. The same experimental setup is used for SFT across all datasets. We use the model after SFT as the base model of RLVC for training. Specific experimental details can be found in Appendix C.
@@ 5.2.1 Evaluation on RECAST-Test
A consistent performance degradation is observed across all models as constraint complexity increases from Level 1 to Level 4. This systematic decline highlights the fundamental challenge LLMs face when processing and satisfying multiple constraints simultaneously. Notably, rule-based constraints present greater difficulties than model-based constraints across all model classes. Among tested models, Gemini-2.5-Pro demonstrates superior constraint handling capabilities with a 39.7539.75% average satisfaction rate.
@@
Models fine-tuned on RECAST-30K demonstrate substantial improvements in constraint satisfaction capabilities. The RECAST-30K-SFT variant achieves a 31.2531.25% average satisfaction rate on Qwen-2.5-7B, significantly outperforming both instruction-tuned models and alternative fine-tuning approaches. This enhancement underscores the effectiveness of our constraint-focused dataset in developing robust instruction-following capabilities, particularly for navigating complex, multi-constraint scenarios.
@@
The application of RLVC further enhances model performance across all difficulty levels. RECAST-30K-RLVC variants consistently outperform their SFT counterparts, achieving a 32.3332.33% average satisfaction rate on Qwen-2.5-7B. This improvement is particularly pronounced in higher difficulty levels (Levels 3-4), where constraint interactions become more complex. These gains confirm that our constraint-specific reinforcement learning approach effectively optimizes for simultaneous satisfaction of multiple requirements, addressing a key challenge in practical instruction following.
@@ 5.3 Ablations and Analysis
To isolate component contributions, we evaluated RECAST variants with specific elements removed. Complete RECAST-30k consistently outperforms all ablated configurations. Instruction-Only Enhancement exhibits the most significant performance degradation (12.2112.21% and 17.3717.37% decreases for Llama and Qwen), revealing that constraint-augmented responses without corresponding response modifications create instruction-response inconsistencies that confuse the model, substantially hampering performance. Response-Only Enhancement performs comparably to Without RECAST, suggesting that constraint specification in instructions drives most performance gains. These results confirm that RECAST’s complete instruction-response synthesis pipeline is essential for optimal constraint satisfaction capabilities.
@@
To analyze the effectiveness of our reinforcement learning approach, we tracked both reward metrics and constraint satisfaction performance throughout the RLVC training process. Figure 5 illustrates the reward progression. This reward optimization corresponds directly to improved constraint satisfaction capabilities. As shown in the evaluation data, the Rule-based Constraint Satisfaction Rate steadily improves from an initial 55% at step 0 to 1111% at step $200$. The reward growth (approximately 55% increase) reflects the inherent difficulty of the constraint satisfaction task. Unlike simpler RL objectives with clearer optimization landscapes, constraint satisfaction requires balancing multiple, potentially competing requirements simultaneously. This complexity creates a challenging reward space where even small improvements represent meaningful learning. These results confirm that our constraint-specific reward design effectively guides model optimization toward improved rule-based constraint handling capabilities, demonstrating the value of targeted reinforcement learning for complex instruction-following tasks.
@@ C.1 Baseline
  • ConiferSun et al. [2024]: Constructs a GPT-4-refined dataset with hierarchical constraint instructions and employs curriculum tuning to enhance complex instruction-following.
  • CrabQi et al. [2024]: Enhances LLMs by back-translating model responses into constraint-rich instructions using a stronger teacher model.
  • I-SHEEPLiang et al. [2024b]: Aligns LLMs from scratch via iterative self-generated instruction–response tuning without external supervision.
  • MUFFINLou et al. [2023]: Diversifies instruction data by scaling input facets per task, yielding more robust multi-faceted alignment.
@@
  • Evol-InstructXu et al. [2024]: Generates complex instructions through iterative rewriting, producing progressively harder instruction–response pairs.
  • SuriPham et al. [2024]: Back-translates long-form texts into multi-constraint prompts, improving structured generation over extended contexts.
  • Tülu 3 Persona IFLambert et al. [2024]: Synthesizes constraint-following data from diverse persona prompts to enhance controllable instruction tuning.
@@ C.2 General Capability Evaluation Benchmark
  • FollowBench Jiang et al. [2024b] comprehensively evaluates instruction-following capability through five types of fine-grained constraints (Content, Situation, Style, Format, and Example) using a multi-level mechanism that incrementally adds constraints to initial instructions. This benchmark systematically measures models’ ability to satisfy multiple constraints simultaneously, with difficulty levels ranging from simple single-constraint instructions to complex multi-constraint scenarios. The evaluation employs metrics including Hard Satisfaction Rate (HSR) for complete constraint satisfaction and Soft Satisfaction Rate (SSR) for partial compliance.
  • IFEval Zhou et al. [2023] assesses models’ ability to follow explicit formatting instructions through strict, verifiable metrics rather than subjective content evaluation. The benchmark focuses on precise adherence to specific requirements such as keyword inclusion, format specifications, and structural constraints. This allows for rigorous, code-based evaluation that eliminates ambiguity in measuring instruction compliance.
  • BBH Suzgun et al. [2023] comprises 23 carefully selected challenging tasks from the BigBench dataset that test diverse cognitive capabilities including multistep arithmetic, algorithmic reasoning, language understanding, and world knowledge. These tasks are chosen for their high difficulty, objective metrics, and strong correlation with human preferences. The benchmark provides insights into models’ reasoning abilities across domains ranging from boolean expression evaluation to sarcasm detection.
  • GPQA Rein et al. [2023] features expert-crafted questions from PhD-level domain specialists in biology, physics, and chemistry designed to be challenging for non-experts yet accessible to specialists. The dataset undergoes rigorous validation to ensure both factual accuracy and appropriate difficulty levels. Access restrictions and gating mechanisms protect against data contamination, maintaining the benchmark’s integrity for evaluating advanced scientific knowledge.
  • MuSR Sprague et al. [2024] tests models’ ability to integrate reasoning with long-context understanding through algorithmically generated complex problems averaging 1,000 words in length. The benchmark includes murder mysteries, object placement puzzles, and team allocation optimization tasks that require maintaining coherent reasoning across extended contexts. Few models achieve above-random performance, making it an effective discriminator of advanced reasoning capabilities.
  • MMLU-PRO Wang et al. [2024b] represents an enhanced version of the original MMLU benchmark, addressing previous limitations through increased difficulty and improved quality control. The benchmark presents 10 answer choices instead of 4, requires deeper reasoning, and underwent expert review to eliminate noisy or unanswerable questions. This refinement results in a more challenging and reliable assessment of multitask language understanding across diverse knowledge domains.